OpenStudy (anonymous):

heeeeeeeeeeeeeeeeeelllllllllllllllllllllppppppppppppppppppppppppppppppppppppppp What is the value of the expression |x| + |y + z| when x = –7, y = 8, and z = –11?

OpenStudy (anonymous):

i got 8-

OpenStudy (anonymous):

Meh, sorry, I'm no good at math...

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

I THINK -9

OpenStudy (anonymous):

or -10

OpenStudy (anonymous):

k

OpenStudy (haleyelizabeth2017):

I think it is 10 :( |-7|+|8-11| 7+|-3| 7+3 10

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

o -10

OpenStudy (anonymous):

-10

OpenStudy (haleyelizabeth2017):

it won't be -10 because it is absolute value!

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

wa

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

It's 10

OpenStudy (haleyelizabeth2017):

You have to do the absolute values before you can add them together.......8-11 is -3 so the abosolute value is 3. the absolute value of -7 is 7 and 7+3 is 10

OpenStudy (anonymous):

so 8 and 9 is wrong?

OpenStudy (haleyelizabeth2017):

yes

OpenStudy (anonymous):

thanks :DDDDDDDDDDDDDDDddd

OpenStudy (haleyelizabeth2017):

no problem :D

OpenStudy (anonymous):

can you help with aother?

OpenStudy (anonymous):

another*

OpenStudy (anonymous):

@haleyelizabeth2017

OpenStudy (anonymous):

What is the solution to the equation x – 0.25 = 1.8?

OpenStudy (anonymous):

What is the solution to the equation x – 0.25 = 1.8?

OpenStudy (anonymous):

x – 0.25 = 1.8

OpenStudy (anonymous):

x- 0.25=1.8

OpenStudy (anonymous):

x-0.25=1.8 add 0.25 to both sides x=1.8+0.25 x=2.05

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

You're welcome

OpenStudy (haleyelizabeth2017):

sorry I couldn't help. I was working on another question and forgot to check back on this one :(

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

does anyone have a roblox?

OpenStudy (anonymous):

nvm